What is vertex of x^2-x-6=0?
Answer by AnlytcPhil(1116)

What is vertex of the parabola whose equation is f(x) = x^2-x-6?
There are two ways to find the vertex of a parabola.
1. Completing the square to get it into the form:
f(x) = a(x-h)² + k and then the vertes is (h,k)
f(x) = 1x² - 1x - 6
Factor the coefficient of x² out of the first two terms:
f(x) = 1(x² - 1x) - 6
Take half the coefficient of x and square it
(1/2)×(1) = 1/2, then (1/2)² = 1/4
Add it then subtract it in the right side of the parentheses:
f(x) = 1(x² - 1x + 1/4 - 1/4) - 6
Change the parentheses to brackets:
f(x) = 1[x² - 1x + 1/4 - 1/4] - 6
Factor the trinomial formed by the first three terms inside the bracket
f(x) = 1[(x - 1/2)(x - 1/2) - 1/4] - 6
Write the factorization as the square of a binomial:
f(x) = 1[(x - 1/2)² - 1/4] - 6
Remove the brackets using the distributive law, leaving the
parentheses intact
f(x) = 1(x - 1/2)² - 1/4 - 6
Combine the two number terms on the right
f(x) = 1(x - 1/2)² - 25/4
Compare to
f(x) = a(x - h)² + k, so
a=1, h=1/2, k=-25/4
So vertex = (h,k) = (1/2,-25/4)
-------
Second way: Use the vertex formula
Vertex formula for f(x) = ax² + bx + c:
Vertex is (h,k) where:
h = -b/(2a)
k = f(h)
f(x) = x² - x - 6
f(x) = 1x² - 1x - 6
a = 1, b = -1, c = -6
h = -b/(2a) = -(-1)/(2·1) = 1/2
k = h² - h - 6 = (1/2)² - (1/2) - 6 = 1/4 - 1/2 - 6 = -25/4
So vertex = (h,k) = (1/2, -25,4)
Edwin